OpenStudy (summersnow8):

it is the final seconds of an ice hockey game between Michigan State and U of M. The Wolverines are down by 1 point. With 20 s left in the game, the Wolverines pull the goalie and have him play as a forward in an attempt to tie the game. The Spartans successfully defend their goal for 9 s. With only .75 s remaining on the game clock, a Spartan shoots the puck on the ice past the skates and sticks of the other players and toward the Wolverines goal. The puck is 26 m from the goal when it leaves the stick with an initial horizontal velocity of 35 m/s. The shot is perfectly directed toward the empty goal, but the ice slows the puck down at a constant rate of 1.50 m/s2 as it slides toward the goal. None of the Wolverines can stop the puck before it reaches the goal. How far does the puck travel when the game clock reaches zero and the horn sounds to end the game?

OpenStudy (anonymous):

Kinematic Motion: X = Xo + Vo * t + 0.5 * a * t^2 Xo = 0 m Vo = 35 m/s t = 0.75 s a = -1.50 m/s^2 X = 0 + (35)(0.75) + (0.5)(-1.50)(0.75)^2 = 18.33 m

OpenStudy (summersnow8):

these are the answer possibilities: 26.25m 25.83m 26.67m -26.6m

OpenStudy (summersnow8):

((35 x .75) + (.5)(-1.50)(.75)squared = 25.83?

OpenStudy (anonymous):

Woops, I made a stupid mistake. My bad.. I punched it in the calculator wrong XD

OpenStudy (anonymous):

I know I wrote it right, that's the correct value to my equation that you got.

OpenStudy (summersnow8):

okay thank you so much for your help